2019/2/5 Commissioning meeting topics
- Assignment of staffs for night work
- Readiness for the X arm commissioning.
- Current status as of 1/22: IR and Green reached X end PD.
1/22: Surface inspection for PR mirrors. Cleaning will be done after 1/31.
1/23: Surface inspection for SR mirrors.
1/24-25: Cleaning HR of SR3, SRM.
1/23: BS is aligned to Y-end PD by IR.
1/27: ITMY is aligned to REFL by IR.
1/27: ITMX is aligned to REFL by IR.
1/27: Beam position of IR at SR3 is confirmed how far from the center.
1/28-29: SR3 is aligned to SR2 center by IR.
1/28: Alignment of SR2 -> SRM by IR.
1/29: Green beam is injected from SR2 AR surface to Y end PD.
1/31: Removing bellows at PR and SR area.
1/31-2/1: Cleaning HR of PR3 and HR, AR of SR2
2/4-8: SRM alignment, SRM -> OMMT by IR -> OMMT -> OSTM by IR, OSTM -> OMC by IR
2/4-8: Alignment of ETMY -> POS by Green
- 2/12-15: Closing door for central area, pumping down
- needs drift monitor for VIS
- Yend (Not yet determined due to the schedule change above)
- 1/25-2/22: Yend Pcal finalization
- 2/25-3/4: Yend WAB
- 3/8~ : Yend pumpdown 3/8~
- 3/11,12: leak test
- Xend (Not yet determined due to the schedule change above)
- 2/22-3/10: Xend Pcal finalization
- 2/11-2/22: Xend NAB installation
- Periscope installation
- 3/11,12 Xend WAB
- 3/19,20, Heatlink etc.
- 3/22 Pumping down
- Leak test
- 3/11~ : Night work start
- 2/18(no delay)-3/8 DRMI commissioning
- 3/11(3weeks delay)-4/5 Y arm commissioning
- 3/25(1week delay)-4/5 FPMI commissioning
http://klog.icrr.u-tokyo.ac.jp/osl/?r=7931: Mirror inspection and cleaning.
JGW-G1909658: new application of photon calibrator for commissioning and observation.
http://klog.icrr.u-tokyo.ac.jp/osl/?r=7942: High power test